# EfficientZero V2: Mastering Discrete and Continuous Control with Limited Data

### v1

@@ 1 Introduction
Previous studies have introduced a range of algorithms aimed at enhancing sample efficiency, including TD-MPC series (Hansen et al., 2022, 2023), EfficientZero (Ye et al., 2021), and the Dreamer series (Hafner et al., 2019, 2021, 2023). Despite these advancements, these algorithms do not consistently attain superior sample efficiency across multiple domains. For example, TD-MPC (Hansen et al., 2022) leverages Model Predictive Path Integral (MPPI) (Rubinstein, 1997) control for planning. The huge computational burden in planning hinders its application in vision-based RL. EfficientZero (Ye et al., 2021) employs the Monte-Carlo Tree Search (MCTS) algorithm, which masters discrete control. However, EfficientZero is unable to handle high-dimensional action spaces, especially in continuous control scenarios. DreamerV3 (Hafner et al., 2023) is a universal algorithm that extends to diverse tasks in a wide range of domains. However, as shown in Fig. 1, DreamerV3 still has noticeable performance gaps to the state-of-the-art (SOTA) algorithms in each domain. Thus, there remains an open question on how to achieve both high performance and sample efficiency across various domains at the same time.
@@ 2.1 Sample Efficient RL
Furthermore, Model-Based Reinforcement Learning (MBRL) has demonstrated high sample efficiency and notable performance in both discrete and continuous control domains. SimPLE (Kaiser et al., 2019), by modeling the environment, predicts future trajectories, thereby achieving commendable performance in Atari games with limited data. TD-MPC (Hansen et al., 2022) utilizes data-driven Model Predictive Control (MPC) (Rubinstein, 1997) with a latent dynamics model and a terminal value function, optimizing trajectories through short-term planning and estimating long-term returns. The subsequent work, TD-MPC2 (Hansen et al., 2023), excels in multi-task environments. In contrast, the TD-MPC series employs MPC to generate imagined latent states for action planning that are 200 times larger than those used in our method, leading to increased computational costs. Dreamer (Hafner et al., 2019), a reinforcement learning agent, develops behaviors from predictions within a compact latent space of a world model. Its latest iteration, Dreamer V3 (Hafner et al., 2023), is a general algorithm that leverages world models and surpasses previous methods across a wide range of domains. It showcases its sample efficiency by learning online and directly in real-world settings (Wu et al., 2022). Although long-horizon planning in Dreamer V3 (Hafner et al., 2023) enhances the quality of the collected data, an imagination horizon of $H=15$ may be excessively long, potentially leading to an accumulation of model errors.
@@ 4.1 Overview
4.2 Training and Planning with a Learned Model
The EZ-V2 algorithm learns a predictive model in a latent space and then performs planning over actions using this model. Specifically, the components of EZ-V2 consist of a representation function, a dynamic function, a policy function, and a value function, which are formulated as follows.
  • Representation Function: $\mathcal{H}:s_{t}=\mathcal{H}(o_{t})$
  • Dynamic Function: $\mathcal{G}:\hat{s}_{t+1},r_{t}=\mathcal{G}(s_{t},\mathcal{A}(a_{t}))$
  • Policy Function: $\mathcal{P}:p_{t}=\mathcal{P}(s_{t})$
  • Value Function : $\mathcal{V}:v_{t}=\mathcal{V}(s_{t})$
Among these components, $o_{t}$ represents the current observation, $s_{t}$ is the current latent state, and $\hat{s}_{t+1}$ is the predicted next state. The representation function learns a compact state representation of the input $o_{t}$. The dynamic function predicts the next state $\hat{s}_{t+1}$ and the reward $r_{t}$. The action embedding block $\mathcal{A}$ encodes the action $a_{t}$ as a latent vector within the dynamic function. The policy function outputs the current policy $p_{t}$, and the value function provides the value estimation $v_{t}$ at the current state. All components are implemented as neural networks, and we provide the details of the architecture for each component in Appendix G.
Figure 2 (A) illustrates the training process of EZ-V2 intuitively. For the dynamic function, we employ a supervised learning method using the true reward $u_{t}$. Furthermore, we introduce temporal consistency (Ye et al., 2021), which strengthens the supervision between the predicted next state $\hat{s}_{t+1}$ and the true next state $s_{t+1}$. Using the learned model, we construct a tree search method for planning actions, specifically a sampling-based Gumbel search. This method can generate the target policy $\pi_{t}$ and target value $z_{t}$ for further supervised learning of the policy and value functions. Simultaneously, it outputs the recommended action $a^{*}_{S}$ for interaction with the environment. The iterative process of interaction and training enhances the accuracy of the dynamic function’s predictions and gradually improves the policy and value functions.
All parameters of the components are trained jointly to match the target policy, value, and reward. Concretely, EZ-V2 is supervised by the following losses:
More details of the training pipeline can be found in Appendix H.
@@ 4.3.1 Target Policy from Tree Search
More specifically, the basic tree search method we adopt is the Gumbel search (Danihelka et al., 2021). This method is recognized for its efficiency in tree searching and its guarantee of policy improvement. At the onset of a search process, the Gumbel search samples $K$ actions using the Gumbel-Top-k trick (Appendix B). It then approaches the root action selection as a bandit problem, aiming to choose the action with the highest Q-value. To evaluate the set of sampled actions, we employ a bandit algorithm known as Sequential Halving (Karnin et al., 2013). However, the Gumbel search primarily investigates planning in discrete action spaces.
@@
Upon completing the sampling-based Gumbel search, we obtain the target policy. We construct two types of target policies. The first is the recommended action $a^{*}_{S}$ obtained from the tree search. In addition to $a^{*}_{S}$, the search also yields $q(s,a)$ for the visited actions. To smooth the target policy, we build the second type as a probability distribution based on the Q-values of root actions (for more details, see Appendix C).
@@ 4.3.2 Learning using Target Policy
Compared with Equation (6), Equation (7) facilitates early exploitation in tasks with a large action dimension, such as the ‘Quadruped walk’ in DM Control (Tassa et al., 2018). We provide an intuitive example to illustrate its advantage in Appendix D.
@@
Normed Median 0.000 1.000 0.144 0.175 0.268 0.415 0.227 1.116 0.490 0.917 1.286
4.4 Search-based Value Estimation
Definition 4.2 (Search-Based Value Estimation).
$\hat{V}_{\text{S}}(s_{0})=\frac{\sum_{n=0}^{N}\hat{V}_{n}(s_{0})}{N}$ (8)
where $H(n)$ denotes the search depth of the $n$-th iteration.
Theorem 4.3 (Search-Based Value Estimation Error).
Define $s_{t},a_{t},r_{t}$ to be the states, actions, and rewards resulting from current policy $\pi$ using true dynamics $\mathcal{G}^{*}$ and reward function $\mathcal{R}^{*}$, starting from $s_{0}\sim\nu$ and similarly define $\hat{s}_{t},\hat{a}_{t},\hat{r_{t}}$ using learned function $\mathcal{G}$. Let reward function $\mathcal{R}$ to be $L_{r}-Lipschitz$ and value function $\mathcal{V}$ as $L_{V}-Lipschitz$. Assume $\epsilon$ as an upper bound of state transition estimation
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]\leq\epsilon^{2}$ (10)
on the compounding error of a tree-search. Then
[…]
The detailed proof can be found in Appendix F.1. SVE possesses several advantageous properties, such as a convergent series coefficient and bounded model errors. Intuitively, the estimation error bound will converge to 0 when the dynamic function approaches optimality, denoted as $\epsilon\to 0$.
Here, $l$ is the horizon for the multi-step TD method. The variable $i_{t}$ denotes the current training step, while $T_{1}$ refers to the initial steps. The term $i_{s}$ indicates the sampling index from the buffer. The buffer size is represented by $|D|$, and $T_{2}$ is the designated threshold for assessing the staleness of data.
@@ 5 Experiment
In this section, we aim to evaluate the general sample efficiency of EZ-V2 on a total of 66 diverse tasks. The tasks include scenarios with low and high-dimensional observation, discrete and continuous action spaces, and dense and sparse rewards. We then present an ablation study on the sampling-based Gumbel search and mixed value target we propose.
@@ 5.3 Ablation Study
Further, we display performance curves for 16 and 8 simulations. As demonstrated in Fig. 3, an increased number of simulations enhances performance in complex tasks such as ‘Quadruped Walk’, and notably, our method with merely 8 simulations surpasses Sample MCTS. This shows that the sampling-based Gumbel search achieves a superior balance between exploration and exploitation, backed by guaranteed policy improvement. Additional results for other tasks are provided in Fig. 9.
@@
This technique also addresses the off-policy issue as the value target is based on the optimal Q-value estimation. Meanwhile, we employ the double Q-head trick. This estimation method is denoted as double Q-value target in Fig. 4. Although the double Q-value target (colored green) also avoids the off-policy issue, the experiments illustrate that our method (colored red) exhibits consistent and robust performance across tasks. This is because our method matches the true value more rapidly by utilizing multi-step predicted rewards during the tree search. Additional curves related to value ablation experiments can be found in Fig. 10.
@@ 6 Conclusion
This paper presents EfficienctZero V2 (EZ-V2), a general framework for sample efficient RL. EZ-V2, which is built upon EfficientZero, extends to master continuous control. Furthermore, EZ-V2 achieves both superior performance and sample efficiency in tasks with visual and low-dimensional inputs. More specifically, EZ-V2 outperforms Dreamer v3 by a large margin across various types of domains, including Atari 100k, Proprio Control, and Vision Control benchmarks. While EZ-V2 presents exciting possibilities, it also comes with certain limitations. Future work should focus on enabling the learned model to handle safety considerations and risks, especially in real-world online learning scenarios like autonomous driving, where stochastic dynamics and interactions with other vehicles pose significant challenges. We encourage collaboration within the research community to tackle these challenges and drive further advancements in the field.
@@ Appendix A Summary of Differences
Appendix B Gumbel-Top-k Trick
$\left(g\in\mathbb{R}^{k}\right)$ $\sim\operatorname{Gumbel}(0)$ (14)
[…]
$\vdots$
Appendix C Calculation of Target Policy
@@
Appendix D Advantage of Simple Policy Loss
@@
Appendix E Details of Search-based Value Estimation (SVE)
E.1 Calculation of SVE
First, the search process expands a tree through $N$ simulations gradually. At each simulation, the agent dives to a leaf and expands a new child. The diving path is easily associated with an $H(n)$-step rollout, which forms an imagined H(n)-step value estimation for the root node. This estimation process will be repeated $N$ times, to get an average estimation, as defined in Definition E.1.
Definition E.1 (Search-Based Value Estimation).
@@ E.1 Calculation of SVE
Appendix F Proof for SVE Error Bound
Theorem F.1 (Search-Based Value Estimation Error).
Define $s_{t},a_{t},r_{t}$ to be the states, actions, and rewards resulting from current policy $\pi$ using true dynamics $\mathcal{G}^{*}$ and reward function $\mathcal{R}^{*}$, starting from $s_{0}\sim\nu$ and similarly define $\hat{s}_{t},\hat{a}_{t},\hat{r_{t}}$ using learned function $\mathcal{G}$. Let reward function $\mathcal{R}$ to be $L_{r}-Lipschitz$ and value function $\mathcal{V}$ as $L_{V}-Lipschitz$. Assume $\epsilon$ as an upper bound of state transition estimation
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]\leq\epsilon^{2}$ (20)
on the compounding error of a tree-search. Then
[…]

### v2

@@ 1 Introduction
Previous studies have introduced a range of algorithms aimed at enhancing sample efficiency, including TD-MPC series (Hansen et al., 2022, 2023), EfficientZero (Ye et al., 2021), and the Dreamer series (Hafner et al., 2019; hafner2021mastering; Hafner et al., 2023). Despite these advancements, these algorithms do not consistently attain superior sample efficiency across multiple domains. For example, TD-MPC (Hansen et al., 2022) leverages Model Predictive Path Integral (MPPI) (Rubinstein, 1997) control for planning. The huge computational burden in planning hinders its application in vision-based RL. EfficientZero (Ye et al., 2021) employs the Monte-Carlo Tree Search (MCTS) algorithm, which masters discrete control. However, EfficientZero is unable to handle high-dimensional action spaces, especially in continuous control scenarios. DreamerV3 (Hafner et al., 2023) is a universal algorithm that extends to diverse tasks in a wide range of domains. However, as shown in Fig. 1, DreamerV3 still has noticeable performance gaps to the state-of-the-art (SOTA) algorithms in each domain. Thus, there remains an open question on how to achieve both high performance and sample efficiency across various domains at the same time.
@@ 2.1 Sample Efficient RL
Furthermore, Model-Based Reinforcement Learning (MBRL) has demonstrated high sample efficiency and notable performance in both discrete and continuous control domains. SimPLE (Kaiser et al., 2019), by modeling the environment, predicts future trajectories, thereby achieving commendable performance in Atari games with limited data. TD-MPC (Hansen et al., 2022) utilizes data-driven Model Predictive Control (MPC) (Rubinstein, 1997) with a latent dynamics model and a terminal value function, optimizing trajectories through short-term planning and estimating long-term returns. The subsequent work, TD-MPC2 (Hansen et al., 2023), excels in multi-task environments. The TD-MPC series employs MPC to generate imagined latent states for action planning. In contrast, our method employs a more efficient action planning module called Sampling-based Gumbel Search, leading to lower computational costs. Dreamer (Hafner et al., 2019), a reinforcement learning agent, develops behaviors from predictions within a compact latent space of a world model. Its latest iteration, Dreamer V3 (Hafner et al., 2023), is a general algorithm that leverages world models and surpasses previous methods across a wide range of domains. It showcases its sample efficiency by learning online and directly in real-world settings (Wu et al., 2022). Although long-horizon planning in Dreamer V3 (Hafner et al., 2023) enhances the quality of the collected data, an imagination horizon of $H=15$ may be excessively long, potentially leading to an accumulation of model errors.
@@ 3.1 Reinforcement Learning
3.2 Gumbel-Top-k Trick
$\left(g\in\mathbb{R}^{k}\right)$ $\sim\operatorname{Gumbel}(0)$ (1)
[…]
$\vdots$
3.3 EfficientZero
3.3.1 Newtork Structure
EfficientZero algorithm learns a predictive model in a latent space and then performs planning over actions using this model. Specifically, the components of EfficientZero consist of a representation function, a dynamic function, a policy function, and a value function, which are formulated as follows.
  • Representation Function: $\mathcal{H}:s_{t}=\mathcal{H}(o_{t})$
  • Dynamic Function: $\mathcal{G}:\hat{s}_{t+1},r_{t}=\mathcal{G}(s_{t},a_{t})$
  • Policy Function: $\mathcal{P}:p_{t}=\mathcal{P}(s_{t})$
  • Value Function : $\mathcal{V}:v_{t}=\mathcal{V}(s_{t})$
Among these components, $o_{t}$ represents the current observation, $s_{t}$ is the current latent state, and $\hat{s}_{t+1}$ is the predicted next state. The representation function learns a compact state representation of the input $o_{t}$. The dynamic function predicts the next state $\hat{s}_{t+1}$ and the reward $r_{t}$. The policy function outputs the current policy $p_{t}$, and the value function provides the value estimation $v_{t}$ at the current state. All components are implemented as neural networks. EZ-V2 uses a similar network structure (see Figure 2), with details of the architecture for each component provided in Appendix G. The main difference is that the action embedding block $\mathcal{A}$ encodes the action $a_{t}$ as a latent vector within the dynamic function.
3.3.2 Training Process
For the dynamic function, EfficientZero employs a supervised learning method using the true reward $u_{t}$. Furthermore, EfficientZero introduces temporal consistency (Ye et al., 2021), which strengthens the supervision between the predicted next state $\hat{s}_{t+1}$ and the true next state $s_{t+1}$. Using the learned model, an MCTS method is used for planning actions. The method can generate the target policy $\pi_{t}$ and target value $z_{t}$ for further supervised learning of the policy and value functions. Simultaneously, it outputs the recommended action $a^{*}_{S}$ for interaction with the environment. The iterative process of interaction and training enhances the accuracy of the dynamic function’s predictions and gradually improves the policy and value functions. EZ-V2 inherits the training process of EfficientZero but replaces the MCTS method with a sampling-based Gumbel search, ensuring policy improvement in continuous action spaces. Figure 2 (A) intuitively illustrates the training process of EZ-V2.
All parameters of the components are trained jointly to match the target policy, value, and reward.
More details of the training pipeline can be found in Appendix H.
@@ 4.2.1 Target Policy from Tree Search
More specifically, the basic tree search method we adopt is the Gumbel search (Danihelka et al., 2021). This method is recognized for its efficiency in tree searching and its guarantee of policy improvement. At the onset of a search process, the Gumbel search samples $K$ actions using the Gumbel-Top-k trick (Section 3.2). It then approaches the root action selection as a bandit problem, aiming to choose the action with the highest Q-value. To evaluate the set of sampled actions, we employ a bandit algorithm known as Sequential Halving (Karnin et al., 2013). However, the Gumbel search primarily investigates planning in discrete action spaces.
@@
Upon completing the sampling-based Gumbel search, we obtain the target policy. We construct two types of target policies. The first is the recommended action $a^{*}_{S}$ obtained from the tree search. In addition to $a^{*}_{S}$, the search also yields $q(s,a)$ for the visited actions. To smooth the target policy, we build the second type as a probability distribution based on the Q-values of root actions (for more details, see Appendix B).
@@ 4.2.2 Learning using Target Policy
Compared with Equation (8), Equation (9) facilitates early exploitation in tasks with a large action dimension, such as the ‘Quadruped walk’ in DM Control (Tassa et al., 2018). We provide an intuitive example to illustrate its advantage in Appendix C.
4.3 Search-based Value Estimation
Definition 4.2 (Search-Based Value Estimation).
$\hat{V}_{\text{S}}(s_{0})=\frac{\sum_{n=0}^{N}\hat{V}_{n}(s_{0})}{N}$ (10)
where $H(n)$ denotes the search depth of the $n$-th iteration.
Corollary 4.3 (Search-Based Value Estimation Error).
Define $s_{t},a_{t},r_{t}$ to be the states, actions, and rewards resulting from current policy $\pi$ using true dynamics $\mathcal{G}^{*}$ and reward function $\mathcal{R}^{*}$, starting from $s_{0}\sim\nu$ and similarly define $\hat{s}_{t},\hat{a}_{t},\hat{r_{t}}$ using learned function $\mathcal{G}$. Let reward function $\mathcal{R}$ to be $L_{r}-Lipschitz$ and value function $\mathcal{V}$ as $L_{V}-Lipschitz$. Assume $\epsilon_{s},\epsilon_{r},\epsilon_{v}$ as upper bounds of state transition, reward, and value estimations respectively. We define the error bounds of each estimation as
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]\leq\epsilon_{s}^{2}$ (12)
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\mathcal{R}(s_{t})-\mathcal{R}^{*}(s_{t})\|^{2}\right]\leq\epsilon_{r}^{2}$ (13)
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\mathcal{V}(s_{t})-\mathcal{V}^{*}(s_{t})\|^{2}\right]\leq\epsilon_{v}^{2}$ (14)
within a tree-search process. Then we have errors
$\text{MSE}_{\nu}(\hat{V}_{\text{S}})$ (15)
$\leq\frac{4}{N^{2}}\sum_{n=0}^{N}\left(\sum_{t=0}^{H(n)}\gamma^{2t}(L_{r}^{2}\epsilon_{s}^{2}+\epsilon_{r}^{2})+\gamma^{2H(n)}(L_{V}^{2}\epsilon_{s}^{2}+\epsilon_{v}^{2})\right)$
The detailed proof can be found in Appendix E.1. SVE possesses several advantageous properties, such as a convergent series coefficient and bounded model errors. Intuitively, the estimation error bound will converge to 0 when the dynamic function approaches optimality, denoted as $\epsilon\to 0$.
Here, $l$ is the horizon for the multi-step TD method. The variable $i_{t}$ denotes the current training step, while $T_{1}$ refers to the initial steps. The term $i_{s}$ indicates the sampling index from the buffer. The buffer size is represented by $|D|$, and $T_{2}$ is the designated threshold for assessing the staleness of data. More details can be found in Appendix F.
@@ 4.3 Search-based Value Estimation
Normed Median 0.000 1.000 0.144 0.175 0.268 0.415 0.227 1.090 0.490 0.917 1.286
@@ 5 Experiment
In this section, we aim to evaluate the general sample efficiency of EZ-V2 on a total of 66 diverse tasks. The tasks include scenarios with low and high-dimensional observation, discrete and continuous action spaces, and dense and sparse rewards. We then present an ablation study on the sampling-based Gumbel search and mixed value target we propose. [^1] The code is released at https://github.com/Shengjiewang-Jason/EfficientZeroV2.
@@ 5.3 Ablation Study
Further, we display performance curves for 16 and 8 simulations. As demonstrated in Fig. 3, an increased number of simulations enhances performance in complex tasks such as ‘Quadruped Walk’, and notably, our method with merely 8 simulations surpasses Sample MCTS. This shows that the sampling-based Gumbel search achieves a superior balance between exploration and exploitation, backed by guaranteed policy improvement. Additional results for other tasks are provided in Appendix J.2.
@@
This technique also addresses the off-policy issue as the value target is based on the optimal Q-value estimation. Meanwhile, we employ the double Q-head trick. This estimation method is denoted as double Q-value target in Fig. 4. Although the double Q-value target (colored green) also avoids the off-policy issue, the experiments illustrate that our method (colored red) exhibits consistent and robust performance across tasks. This is because our method matches the true value more rapidly by utilizing multi-step predicted rewards during the tree search. Additional curves related to value ablation experiments can be found in Appendix J.2.
Furthermore, practical comparisons between the TD-MPC2 and EZ-V2 algorithms in terms of computational load are provided in Appendix J.3.
@@ 6 Conclusion
This paper presents EfficienctZero V2 (EZ-V2), a general framework for sample efficient RL. EZ-V2, which is built upon EfficientZero, extends to master continuous control. Furthermore, EZ-V2 achieves both superior performance and sample efficiency in tasks with visual and low-dimensional inputs. More specifically, EZ-V2 outperforms Dreamer V3 by a large margin across various types of domains, including Atari 100k, Proprio Control, and Vision Control benchmarks. Moreover, we evaluate the performance of EZ-V2 without conditions of limited data. EZ-V2 outperforms or is comparable to baselines with more interaction data, though the performance gap between EZ-V2 and other algorithms narrows as the amount of interaction data increases. Due to the high sample efficiency, we will broaden our evaluation of EZ-V2 across a wider variety of benchmarks, especially in real-world tasks like robotic manipulation. The superior sample efficiency of EZ-V2 holds great promise for enhancing online learning in real-world robotic scenarios.
Acknowledgements
This work is supported by the Ministry of Science and Technology of the People’s Republic of China, the 2030 Innovation Megaprojects “Program on New Generation Artificial Intelligence” (Grant No. 2021AAA0150000). This work is also supported by the National Key R&D Program of China (2022ZD0161700).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Summary of Differences
Appendix B Calculation of Target Policy
@@
Appendix C Advantage of Simple Policy Loss
@@
Appendix D Details of Search-based Value Estimation (SVE)
D.1 Calculation of SVE
First, the search process expands a tree through $N$ simulations gradually. At each simulation, the agent dives to a leaf and expands a new child. The diving path is easily associated with an $H(n)$-step rollout, which forms an imagined H(n)-step value estimation for the root node. This estimation process will be repeated $N$ times, to get an average estimation, as defined in Definition D.1.
Definition D.1 (Search-Based Value Estimation).
@@ D.1 Calculation of SVE
Appendix E Proof for SVE Error Bound
Corollary E.1 (Search-Based Value Estimation Error).
Define $s_{t},a_{t},r_{t}$ to be the states, actions, and rewards resulting from current policy $\pi$ using true dynamics $\mathcal{G}^{*}$ and reward function $\mathcal{R}^{*}$, starting from $s_{0}\sim\nu$ and similarly define $\hat{s}_{t},\hat{a}_{t},\hat{r_{t}}$ using learned function $\mathcal{G}$. Let reward function $\mathcal{R}$ to be $L_{r}-Lipschitz$ and value function $\mathcal{V}$ as $L_{V}-Lipschitz$. Assume $\epsilon_{s},\epsilon_{r},\epsilon_{v}$ as upper bounds of state transition, reward, and value estimations respectively. We define the error bounds of each estimation as
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]\leq\epsilon_{s}^{2}$ (22)
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\mathcal{R}(s_{t})-\mathcal{R}^{*}(s_{t})\|^{2}\right]\leq\epsilon_{r}^{2}$ (23)
$\max_{n\in[N],t\in[H(n)]}\mathbb{E}\left[\|\mathcal{V}(s_{t})-\mathcal{V}^{*}(s_{t})\|^{2}\right]\leq\epsilon_{v}^{2}$ (24)
within a tree-search process. Then we have errors
$\text{MSE}_{\nu}(\hat{V}_{\text{S}})\leq\frac{4}{N^{2}}\sum_{n=0}^{N}\left(\sum_{t=0}^{H(n)}\gamma^{2t}(L_{r}^{2}\epsilon_{s}^{2}+\epsilon_{r}^{2})+\gamma^{2H(n)}(L_{V}^{2}\epsilon_{s}^{2}+\epsilon_{v}^{2})\right)$ (25)